# CooPre: Cooperative Pretraining for V2X Cooperative Perception

### v1

@@ IV-A Experiment Setting
Datasets. We evaluate our method on three datasets, namely V2X-Real[1], V2V4Real[8], and OPV2V[21]. V2X-Real is a large-scale, real-world V2X dataset that encompasses all V2X collaboration modes, including vehicle-centric (VC), infrastructure-centric (IC), vehicle-to-vehicle (V2V), and infra-to-infra (I2I). LiDAR sensors for this dataset were embodied with different sensor configurations and deployed in both intersection and corridor scenarios. Note that except for the V2V subset in the V2X-Real dataset, the rest of the subsets all exhibit such LiDAR sensor heterogeneity property. We also examine the effectiveness of our pretraining method on V2V4Real and OPV2V, two well-established benchmarks for real-world and simulated V2V cooperative perception.
Evaluation Metrics. Following previous evaluation protocols [1, 21, 8], we adopt the Average Precisions (AP) at the specified Intersection-over-Union (IoU) threshold as the metric to evaluate the detection performance. For V2X-Real [1] dataset, we evaluate the performance of different subclasses (i.e. vehicle, pedestrian, and truck) and a final mean average precision (mAP) at the threshold of 0.3 and 0.5. For V2V4Real [8] dataset, we evaluate the performance of the vehicle class at the threshold of 0.5 and 0.7 in both synchronized and asynchronized modes. For OPV2V [21] dataset, we evaluate the performance of the vehicle class at the threshold of 0.5 and 0.7 at their two separate test sets.
Implementation Details. We evaluate our pretraining method using SECOND backbone[30]. Thus to have a fair comparison, we also train SECOND baseline method in V2X-Real[1] and V2V4Real[8]. All experiments are conducted in one Nvidia A6000 GPU. We employ AdamW[49] optimizer with a weight decay of $1\times 10^{-2}$ to optimize our models. During the pretraining stage, we train the model with a batch size of 4 for 15 epochs using a learning rate of 0.002, and we decay the learning rate with a cosine annealing[50]. We use a masking ratio of 0.7 in our main experiments and a fixed predicted point cloud number of 20. During the fine-tuning stage, the optimization process is identical to the train-from-scratch baselines. We finetune the model for 40 epochs in OPV2V dataset, 60 epochs in V2V4Real dataset, and 20 epochs in V2X-Real dataset. We also add normal point cloud data augmentations for all experiments, including scaling, rotation, and flip[28].
TABLE IV: Results of cross-domain transferability of our method. Results are displayed in terms of AP0.5 of the Car category. Green denotes the improvements over the train-from-scratch baseline.
@@
Figure 5: Data efficiency experiment. We evaluate the performance boost using CooPre under different ratios of finetuning data on V2X-Real dataset.
Figure 6: Comparison of CooPre across different cooperative fusion strategies. We assess the performance enhancement provided by CooPre under various fusion strategies on the V2X-Real VC and V2V datasets.
Figure 7: Robustness assessment on localization errors and time delay. Our method shows strong robustness against localization errors and time delay compared to the train-from-scratch baseline on V2V4Real dataset.
TABLE V: Effectiveness of cooperative agents during pretraining. Results are displayed in terms of mAP0.5.
Dataset Pretraining Mode Overall 0-30m 30-50m 50-100m
@@
Ego pretraining 51.1 64.5 55.7 39.4
V2X pretraining 52.2 64.9 55.9 42.1
@@
Ego pretraining 50.5 65.4 54.6 38.7
V2V pretraining 53.1 66.1 55.4 42.7
@@ IV-B Main Results
In Table I, we present the experimental results across all collaboration modes in V2X-Real[1] dataset. In the VC and V2V settings, CooPre substantially outperforms the train-from-scratch baselines across all classification categories. The improvements are particularly notable in the Car and Truck categories, with a 4.4 mAP0.5 increase in the Car category and a 6.3 mAP0.5 increase in the Truck category in the VC setting, and a 4.8 mAP0.5 increase in the Car category and a 5.3 mAP0.5 increase in the Truck category in the V2V setting. This aligns with our pretraining design, which enables the model to learn more 3D geometrical and topological information beforehand, resulting in more accurate bounding box detection results with higher thresholds. On the other hand, the pretraining shows an incremental effect on detecting pedestrians. We attribute this to two factors: 1) pedestrians are small-scale in nature and thus receive fewer LiDAR features than larger objects, and 2) pedestrians are non-rigid-body objects, making it more challenging for the model to learn their 3D features compared to rigid-body counterparts such as cars and trucks. On the infrastructure side, our pretraining method shows improvements for the Car and Truck categories, but we observe a subtle decrease in the Pedestrian category. We attribute this to the small-scale and non-rigid-body nature of pedestrians, which might be negatively affected by asynchronized results transmitted from connected vehicles, such as pose and localization errors[13, 14]. This asynchronization can alter the perception accuracy of static infrastructure observers.
We show the generalizability of our method in another real-world dataset V2V4Real[8] and a simulation dataset OPV2V[21], as shown in Table II and Table III, respectively. For the V2V4Real dataset, after we pretrain and finetune the model, we test it under synchronized and asynchronized modes for a fair comparison. The improvement is substantial in both testing settings. For the OPV2V dataset, we outperform the baseline by a large margin across all test sets. These results demonstrate the generalizability of our pretraining method across different domains.
@@ IV-C Data Efficiency
In this section, we investigate the benefits of our pretraining method in scenarios with limited labeled data. Specifically, we randomly sample 20%, 50%, and 80% of the training dataset and train the models with these annotated subsets. For our method, we pretrain the model on the entire training set and then finetune it on each sampled subset. As shown in Fig.5, our method outperforms train-from-scratch baselines across all settings. Notably, the performance gain of CooPre increases as the percentage of data decreases. Additionally, CooPre provides crucial guidance when collaboration involves different sensor configurations. For instance, in the V2X-Real VC dataset, the baseline method struggles to learn a good representation with limited annotations and the performance drops dramatically due to data heterogeneity issues. With our pretraining method, the model learns meaningful prior knowledge of different sensor distributions, leading to substantial improvements even when finetuning with less labeled data. These findings demonstrate the effectiveness of our method in data scarcity scenarios.
@@ IV-D Cross-domain Transferability
We evaluate the cross-dataset transferability of our pretrained 3D encoder by finetuning it on another dataset. To isolate the effects of data heterogeneity, we investigate the performance across two real-world V2V datasets (V2X-Real V2V and V2V4Real), which have similar sensor configurations. As shown in Table IV, pretraining on a different domain improves performance compared to the train-from-scratch baseline. However, it performs worse than pretraining on the source domain due to the domain gap issue. We also conduct an experiment where we combine the V2X-Real V2V and V2V4Real datasets to create a large pretraining corpus. While this approach improves performance over pretraining on the source domain, the gains are less significant. This difference could be attributed to the scenario differences between the V2X-Real V2V and V2V4Real datasets. The former primarily focuses on intersection scenarios, whereas the latter includes a broader range of corridor scenarios in its training set.
IV-E Ablation Studies
@@
Effectiveness of cooperative agents during pretraining. To assess the impact of collaboration during pretraining, we compare the performance of our multi-agent pretraining method with the ego-agent pretraining method[18]. As shown in Table V, multi-agent cooperative pretraining leads to further improvements in model performance.
Extent of improvements in different perception ranges. We also explore the enhancements across various perception ranges. As shown in TableV and TableII, the most substantial improvements are observed in middle and long ranges compared to the baselines. Given that our multi-agent pretraining method achieves a significantly larger perception field compared to the ego-agent pretraining method, our approach exhibits greater robustness in handling long-range perception and occlusions, as illustrated in Fig.4.
@@
Robustness assessment. Following [2], we evaluate our method’s robustness towards localization error and time delay compared to the baseline method on the V2V4Real dataset, as illustrated in Fig.7. Our method is less sensitive to localization errors and shows its strong robustness against time delay.
Masking Ratio. Table VI demonstrates the effect of the masking ratio within the range of 0.6 to 0.8. With a masking ratio of 0.7, the pretraining framework achieves the best performance.
@@
Dataset Mask Ratio mAP0.3 mAP0.5
V2X-Real VC 0.6/0.7/0.8 59.1/60.1/59.0 51.7/52.2/51.0
V2X-Real V2V 0.6/0.7/0.8 57.6/59.5/58.0 51.4/53.1/51.6
@@ V Conclusion
We introduce a multi-agent pretraining framework that prompts the representation to learn a holistic prior knowledge of the 3D environment before performing the perception task. The framework explores the intrinsic geometrical and topological information of scenarios and sensor distributions. Extensive experiments on representative datasets demonstrate the efficacy of the method as it outperforms the previous state-of-the-art methods in all V2X settings. Furthermore, we demonstrate this framework’s strong generalizability, cross-domain adaptability, and data efficiency in cooperative perception. Future work includes extending this self-supervised cooperative pretraining paradigm to the field of cooperative perception and prediction task.

### v2

@@ IV-A Experiment Setting
Datasets. We evaluate CooPre on three datasets, namely V2X-Real [1], V2V4Real [8], and OPV2V [18]. V2X-Real is a large-scale, real-world V2X dataset that encompasses all V2X collaboration modes, including vehicle-centric (VC), infrastructure-centric (IC), vehicle-to-vehicle (V2V), and infra-to-infra (I2I). LiDAR sensors for this dataset were embodied with different sensor configurations and deployed in both intersection and corridor scenarios. Note that except for the V2V subset in the V2X-Real dataset, the rest of the subsets all exhibit such LiDAR sensor heterogeneity property. We also examine the effectiveness of CooPre on V2V4Real and OPV2V, two well-established benchmarks for real-world and simulated V2V cooperative perception.
Evaluation Metrics. Following previous evaluation protocols [1, 18, 8], we adopt the Average Precisions (AP) at the specified Intersection-over-Union (IoU) threshold as the metric to evaluate the detection performance. For V2X-Real [1] dataset, we evaluate the performance of different subclasses (i.e. car, pedestrian, and truck) and a final mean average precision (mAP) at the threshold of 0.3 and 0.5. For V2V4Real [8] dataset, we evaluate the performance of the vehicle class at the threshold of 0.5 and 0.7 in both synchronized and asynchronized modes. For OPV2V [18] dataset, we evaluate the performance of the vehicle class at the threshold of 0.5 and 0.7 at their two separate test sets. Green denotes the improvements over the corresponding backbone for all experiments.
Implementation Details. We evaluate CooPre using SECOND backbone [30] since it obtains higher performance and faster training speed in the real-world dataset [1, 8]. Note that CooPre could also be generalized to other 3D encoders such as PointPillar [28], as shown in Table VII. All experiments are conducted in one Nvidia A6000 GPU. We employ AdamW [51] optimizer with a weight decay of $1\times 10^{-2}$ to optimize our models. During the pretraining stage, we train the model with a batch size of 4 for 15 epochs using a learning rate of 0.002, and we decay the learning rate with a cosine annealing [52]. We use a masking ratio of 0.7 in our main experiments and a fixed predicted point cloud number of 20. During the fine-tuning stage, the optimization process is identical to the train-from-scratch baselines. We finetune the model for 40 epochs in OPV2V dataset, 60 epochs in V2V4Real dataset, and 20 epochs in V2X-Real dataset. We also add normal point cloud data augmentations for all finetuning experiments, including scaling, rotation, and flip [28]. Note that both pretraining and finetuning stages are conducted within the same dataset.
Figure 5: Data efficiency experiment. We evaluate the performance boost using CooPre under different ratios of finetuning data on V2X-Real dataset with AttFuse [18] as backbone.
TABLE IV: Results of cross-domain transferability of our method (Metric: AP0.5 for Car Category).
@@
Figure 6: Comparison of CooPre across different cooperative fusion strategies. We assess the performance enhancement provided by CooPre under various fusion strategies on the V2X-Real VC and V2V datasets with AttFuse [18] as backbone.
Figure 7: Robustness assessment on localization errors and time delay. Our method shows strong robustness against localization errors and time delay compared to the train-from-scratch baseline on V2V4Real dataset with AttFuse [18] as backbone.
TABLE V: Comparison with prior works (Metric: mAP0.5).
Dataset Pretraining Method Overall 0-30m 30-50m 50-100m
@@
BEVContrast [21] 48.8 61.2 52.5 39.2
Random Masking 50.9 63.9 55.4 40.3
CORE [50] 51.6 64.7 55.4 40.7
STAR [53] 50.0 63.9 55.6 38.4
BEV-MAE [20] 51.1 64.5 55.7 39.4
Ours 52.2 64.9 55.9 42.1
@@
BEVContrast [21] 46.3 63.1 54.1 32.9
Random Masking 50.9 65.5 54.7 40.4
CORE [50] 51.9 65.9 55.0 41.8
STAR [53] 51.2 66.0 54.4 40.6
BEV-MAE [20] 50.5 65.4 54.6 38.7
Ours 53.1 66.1 55.4 42.7
@@ IV-B Main Results
In Table I, we present the experimental results across all collaboration modes in V2X-Real [1] dataset. In the VC and V2V settings, CooPre substantially outperforms the train-from-scratch baselines across all classification categories. Using AttFuse [18] as an example, improvements are especially evident in the Car and Truck categories, with AP0.5 gains of 4.4 and 6.3 in the VC setting, and 4.8 and 5.3 in the V2V setting, respectively. This aligns with our pretraining design, which enables the model to learn more 3D geometrical and topological information beforehand, resulting in more accurate bounding box detection results under higher thresholds. On the other hand, our pretraining method shows an incremental effect on detecting pedestrians. We attribute this to two factors: 1) pedestrians are small-scale in nature and thus receive fewer LiDAR features than larger objects, and 2) pedestrians are non-rigid-body objects, making it more challenging for the model to learn their 3D features compared to rigid-body counterparts such as cars and trucks. On the infrastructure side, our pretraining method shows improvements for the Car and Truck categories, but we observe a subtle decrease in the Pedestrian category. We attribute this to the small-scale and non-rigid-body nature of pedestrians, which might be negatively affected by asynchronized results transmitted from connected vehicles, such as pose and localization errors [13, 14]. This asynchronization can alter the perception accuracy of static infrastructure observers. Such observation is also shown in the results of other cooperative methods [19, 2].
We show the generalizability of CooPre in the V2V4Real [8] and OPV2V [18] datasets, as shown in Table II and Table III, respectively. For the V2V4Real dataset, after we pretrain and finetune the model, we test it under synchronized and asynchronized modes for a fair comparison following the original paper settings. The improvement is substantial in both testing settings. For the OPV2V dataset, we outperform the baseline by a large margin across all test sets. These results demonstrate the generalizability of CooPre across different domains with different fusion methods.
@@ IV-C Data Efficiency
In this section, we investigate the benefits of our pretraining method in scenarios with limited labeled data using AttFuse [18] as backbone. Specifically, we randomly sample 20%, 50%, and 80% of the training dataset and train the models with these annotated subsets. For our method, we pretrain the model on the entire training set and then finetune it on each sampled subset. As shown in Fig. 5, our method outperforms train-from-scratch baselines across all settings. Notably, the performance gain of CooPre increases as the percentage of data decreases. Additionally, CooPre provides crucial guidance when collaboration involves different sensor configurations. For instance, in the V2X-Real VC dataset, the baseline method struggles to learn a good representation with limited annotations and the performance drops dramatically due to data heterogeneity issues. With our pretraining method, the model learns meaningful prior knowledge of different sensor distributions, leading to substantial improvements even when finetuning with less labeled data. These findings demonstrate the effectiveness of our method in data scarcity scenarios.
@@ IV-D Cross-domain Transferability
We evaluate the cross-dataset transferability of our pretrained 3D encoder by finetuning it on another dataset using AttFuse [18] backbone. Since V2V4Real [8] does not support multi-class classification, we only examine the perception results on car category which is shared by both domains. To isolate the effects of data heterogeneity, we investigate the performance across two real-world V2V datasets (V2X-Real V2V and V2V4Real), which have similar sensor configurations. As shown in Table IV, pretraining on a different domain improves performance compared to the train-from-scratch baseline. However, it performs worse than pretraining on the source domain due to the domain gap issue. We also conduct an experiment where we combine the V2X-Real V2V and V2V4Real datasets to create a large pretraining corpus. While this approach improves performance over pretraining on the source domain, the gains are less significant. This difference could be attributed to the scenario differences between the V2X-Real V2V and V2V4Real datasets. The former primarily focuses on intersection scenarios, whereas the latter includes a broader range of corridor scenarios in its training set.
IV-E Ablation Studies and Discussion
@@
Comparison with other pretraining methods. The comparison of pretraining methods is carried out in three aspects using AttFuse [18] as backbone. First, to evaluate reconstruction targets in a multi-agent pretraining setting, we adapt the reconstruction loss function from [50, 53] to align with our pretraining framework. Our approach demonstrates superior performance. Second, to analyze the impact of collaboration during pretraining, we compare our multi-agent pretraining method with the ego-agent pretraining approach from [20, 21]. As shown in Table V, multi-agent pretraining results in further performance gains. Additionally, both single-agent and multi-agent pretraining outperform the train-from-scratch approach. Lastly, we note variability in [21], a contrastive learning based single-agent pretraining method, which may struggle to establish robust priors for multi-agent BEV representation compared to reconstruction-based pretraining methods.
Extent of improvements in different perception ranges. We also explore the enhancements across various perception ranges. As shown in Table V and Table II, the most substantial improvements are observed in middle and long ranges compared to the baselines. Since our multi-agent pretraining method enables a significantly larger perception field and provides supervision signals for feature learning on cooperative agents, it demonstrates greater robustness in handling long-range perception and occlusions compared to single-agent pretraining methods, as illustrated in Fig. 4.
@@
Robustness assessment. Following [2], we evaluate our method’s robustness towards localization error and time delay on the V2V4Real dataset, as illustrated in Fig. 7. Our method is less sensitive to localization errors and shows strong robustness against time delay.
Masking Ratio. Table VI demonstrates the effect of the masking ratio within the range of 0.6 to 0.8. With a masking ratio of 0.7, CooPre achieves the best performance.
Effectiveness on different 3D encoders. Table VII shows CooPre consistently improves the detection accuracy with different 3D encoders [30, 28].
@@
Dataset Method Mask Ratio mAP0.3 mAP0.5
V2X-Real VC AttFuse [18] 0.6/0.7/0.8 59.1/60.1/59.0 51.7/52.2/51.0
TABLE VII: Ablation on different 3D encoders in terms of mAP0.3/0.5.
Dataset Method Second PointPillar
Original W/ CooPre Original W/ CooPre
V2X-Real VC AttFuse [18] 56.1/48.5 60.1/52.2 42.5/33.8 44.0/35.2
@@ V Conclusion
We introduce CooPre, a multi-agent pretraining framework that prompts the representation to learn a holistic prior knowledge of the 3D environment before performing the perception task. The framework explores the intrinsic geometrical and topological information of scenarios and sensor distributions. Extensive experiments on representative datasets demonstrate the efficacy of the method as it outperforms the previous state-of-the-art methods in all V2X settings. Furthermore, we demonstrate this framework’s strong generalizability, cross-domain adaptability, and data efficiency in cooperative perception. Future work includes extending this self-supervised cooperative pretraining paradigm to the field of cooperative perception and prediction task.
ACKNOWLEDGMENT
This paper is partially supported by USDOT/FHWA Mobility Center of Excellence, National Science Foundation # 2346267 POSE: Phase II: DriveX.